solve the logarithmic equation. be sure to reject any value that is not in the domain of the original logarithmic expression. Give the exact answer.
log2(x-1) = 1 + log2 (x-5)

Answer by oscargut(891)

log2(x-1) = 1 + log2 (x-5)
log2(x-1) - log2 (x-5)= 1
log2((x-1)/(x-5))=1 (because log(a)-log(b)=log(a/b))
x-1/x-5=2
x-1=2x-10
-x+9=0
x=9
checking:
log2(8)=1+log2(4)
3=1+2
Answer: x=9